OpenStudy (anonymous):

1. What is the slope of the line passing through (0,3) and (4,5)? 2. Write the slope-intercept form of the equation of the line that passes through (4,2) with a slope of 3. 3. Which of the following is the equation of the vertical line that passes through (4,6)?

OpenStudy (anonymous):

for 1 use the formula to find the slope, m=y2-y1/x2-x1 for 2 use the slope and point given with the formula y-y1=m(x-x1), put it in the form y=mx+b not sure about 3

OpenStudy (anonymous):

@Bunki<3 so is the answer to number 1, -0.5?

OpenStudy (anonymous):

@Bunki<3 or it is -2?

OpenStudy (anonymous):

(5-3)/(4-0)